OpenStudy (anonymous):

5=1+5sin(pi/2)(x+0.7) I'm supposed to find the first three positive values. I know how to do this with degrees, but I'm stuck when it comes to doing it in radians.

OpenStudy (freckles):

\[5=1+5 \cdot \sin(\frac{\pi}{2}) \cdot (x+\frac{7}{10}) ?\]

OpenStudy (anonymous):

\[5=1+5\sin \frac{ \pi }{ 2 } (x+0.7)\]

OpenStudy (freckles):

what is that sin function operating on?

OpenStudy (freckles):

just the pi/2

OpenStudy (freckles):

or pi/2(x+7/10)?

OpenStudy (freckles):

\[5=1+5 \cdot \sin(\frac{\pi}{2}(x+\frac{7}{10}) ) ? \\ \text{ I think you mean this }\]

OpenStudy (freckles):

First isolate the part that is \[\sin(\frac{\pi}{2} \cdot (x+\frac{7}{10}) ) \]

OpenStudy (anonymous):

No the one I typed out in my response is the one I have in my textbook. I've gotten as far as \[2\pi n \pm 0.93 = \frac{ \pi }{ 2 }(x+0.7)\]

OpenStudy (anonymous):

I'm supposed to divide both sides by \[\frac{ \pi }{ 2 }\] but I keep getting the wrong answer after that.

OpenStudy (freckles):

try multiplying by the reciprocal of pi/2 on both sides instead of dividing both sides by pi/2 looks at like that might make it less complicated for you

OpenStudy (anonymous):

Is that supposed to give me \[4\pi \pm 1.86\]?

OpenStudy (freckles):

\[\frac{2}{\pi}(2 \pi n \pm 0.93)=x+0.7\]

OpenStudy (freckles):

and no

OpenStudy (anonymous):

\[4\pi \pm 0.592\]? because I got that and tried to complete the equation from there and I still got the wrong answer. Not really sure what I'm doing incorrectly.

OpenStudy (freckles):

well i just said no

OpenStudy (freckles):

\[\frac{2}{\pi}(2 \pi n \pm 0.93)=x+0.7 \\ \text{ distribute } \\ \frac{2 \cdot 2 \pi n }{ \pi } \pm \frac{2}{\pi} 0.93=x+0.7\]

OpenStudy (freckles):

the pi's would cancel in the first term

OpenStudy (freckles):

also do you know that sin(-0.93) isn't 4/5

OpenStudy (anonymous):

I was supposed to take the arcsin of 4/5 which is 0.93

OpenStudy (freckles):

right sin(.93) is 4/5 but sin(-.93) isn't 4/5

OpenStudy (freckles):

anyways .93 is between 0 and pi/2 so another number that occurs between 0 and 2pi pi-.93 is another that occurs between 0 and 2pi such that the sin value is 4/5 I know this from looking at the unit circle

OpenStudy (freckles):

So this means you would have sin(.93+2pi*n)=4/5 or sin((pi-.93)+2pi*n)=4/5

OpenStudy (freckles):

where n is an integer

OpenStudy (freckles):

But anyways you still need to solve \[.93+2 \pi \cdot n =\frac{\pi}{2}(x+.7) \\ \text{ and } \pi-.93+2\pi \cdot n=\frac{\pi}{2}(x+.7)\]